Macroscale Crystallization pre lab calculations:

1. Calculate how much 95% ethyl alcohol will be required to dissolve 0.75 g of sulfanilamide at 78 degrees Celsius.

2. Using the volume of solvent calculated in step 1, calculate how much sulfanilamide will remain dissolved in the mother liquor after the mixture is cooled to 0 degrees Celsius.
Answer #1
  1. Solubility of sulfanilamide in 95% ethanol at 78 degrees Celsius is 210 mg/ml.

We need to dissolve 0.75 g = 750 mg of sulfanilamide

210 mg dissolves in 1 ml

750 mg will dissolve in = (1/210)×750 = 3.57 ml

3.57 ml of  95% ethyl alcohol will be required to dissolve 0.75 g of sulfanilamide at 78 degrees Celsius.

  1. Solubility of sulfanilamide in 95% ethanol at 0 degrees Celsius is 14 mg/ml.

We need to calculate how much sulfanilamide stll remain dissolve at 0 degrees Celsius out of 0.75 g = 750 mg of sulfanilamide.

1 ml contains 14 mg sulfanilamide at 0 degrees Celsius

3.57 ml will contain = 14 × 3.57 = 50 mg = 0.05 g

Using the volume of solvent calculated in step 1, 0.05 g sulfanilamide will remain dissolved in the mother liquor after the mixture is cooled to 0 degrees Celsius.